OpenStudy (anonymous):

What did I do wrong in this problem? (Finding limits with trigonometric functions)

OpenStudy (anonymous):

I know my finaly answer is wrong, but I can't figure out what I did wrong. I know there are other ways to solve it, but what exactly is invalid about the steps i took.

OpenStudy (misty1212):

HI!

OpenStudy (anonymous):

hey :D

OpenStudy (misty1212):

before we begin, is it clear that the answer is \(\frac{2}{4}=\frac{1}{2}\) ? they are after all the only numbers you see

OpenStudy (anonymous):

Yeah, I'm pretty sure looking at the graph the final answer should be 0.5. However the way I did it, my answer comes to 0.

OpenStudy (misty1212):

this is what you know:\[\lim_{x\to 0}\frac{\sin(x)}{x}=1\]

OpenStudy (misty1212):

and therefore \[\lim_{x\to 0}\frac{\sin(ax)}{ax}=1\] as well since they are the same thing

OpenStudy (misty1212):

so the gimmick here is to write \[\lim_{x\to 0}\frac{\sin(ax)}{\sin(bx)}=\lim_{x\to 0}\frac{\sin(ax)}{ax}\frac{bx}{\sin(bx)}\times \frac{ax}{bx}\]

OpenStudy (misty1212):

you get \[1\times 1\times \frac{a}{b}\]

OpenStudy (misty1212):

which in your specific example is \(\frac{2}{4}\)

OpenStudy (anonymous):

Oh I see. I think I factored out the limits in the wrong way. Alright thanks for the answer

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]